# Ultrafast Band Structure Dynamics in Bulk 1T-VSe

### v1

@@ Abstract
Complex materials encompassing different phases of matter can display new photoinduced metastable states differing from those attainable under equilibrium conditions. These states can be realized when energy is injected in the material following a non-equilibrium pathway, unbalancing the unperturbed energy landscape of the material. Guided by the fact that photoemission experiments allow for detailed insights in the electronic band structure of ordered systems, here we study bulk 1T-VSe in its metallic and charge-density-wave phase by time- and angle-resolved photoelectron spectroscopy. After near-infrared optical excitation, the system shows a net increase of the density of states in the energy range of the valence bands, in the vicinity of the Fermi level, lasting for several picoseconds. We discuss possible origins as band shifts or correlation effects on the basis of a band structure analysis. Our results uncover the possibility of altering the electronic band structure of bulk 1T-VSe for low excitation fluences, contributing to the understanding of light-induced electronic states.
@@
Layered transition-metal dichalcogenides (TMDCs) are attracting particular attention since they exhibit a variety of physical properties such as spin- and valley-coupled two-dimensional superconductivity, exciton Hall effect and exotic charge-density-wave (CDW) transitions.[1]
@@
Path breaking works on light induced CDWs in compunds as VO[2, 3] and LaTe [4] triggered the search for similar effects in other types of materials, as TMDCs.
@@
Photoexcitation can drive TMDC-CDW systems in metastable states surviving for several picoseconds, as discussed for the examples of 1-TiSe[5, 6], 1-TaSe[7] and 1-VSe[8].
@@
Origins for these states are attributed to phonon bottle-neck effects and optically induced phase transitions from the CDW phase into a metallic phase.
@@
Duan et al. discovered that for TiSe an optically induced metastable metallic phase can be induced far below the CDW phase transition temperature, showing different properties than the equilibrium normal phase.[6]
@@
As the origin the optical triggering of a halted motion of the atoms via coherent electron-phonon coupling processes was discussed.
@@
In the case of 1T-VSe instead the phonon bottle-neck effects were discussed after optical excitation in its metallic phase.[8]
@@
Recent interest has been drawn to the monolayer limit of VSe due to the enhanced CDW transition temperature and enlarged band gap.[9, 10, 11, 12]
@@
Biswas et al.[9] report on ultrafast light induced insulator-metal transition in 2D VSe.
@@
However, there are still open questions regarding the bulk properties of the material to be answered to gain a profound understanding of the physical mechanisms that determine the materials properties and phase transitions.
@@
With this work we contribute to the understanding of the optically induced metastable state of 1T-VSe by performing time- and angle-resolved photoemission (tr-ARPES) experiments.
@@
We excite the material with 1.55 eV pump pulses and probe at its point with 200 fs time and 50 meV energy resolution.
@@
Comparing results acquired at three different sample temperatures (well above, around and well below the CDW transition temperature of T110 K[13]), we also address the influence of the initial phase of the system on the observed metastable modification.
@@
ARPES gives access to the valence band structure of a material, which in the case of 1T-VSe is given by the vanadium 3d (V) and selenium 4p (Se) bands.[14, 15, 8]
@@
Beyond the equilibrium band structure determination, tr-ARPES allows studying the material response to an optical stimulus that drives the system far from equilibrium.
@@
This technique allows to study band-specific carrier dynamics which reveal e.g. information about the electron-electron and electron-phonon interactions in the system under investigation.
Moreover, Tr-ARPES has been shown to be an adequate technique to study light-induced metastable states in the class of TMDC materials and beyond.[8, 16, 6, 5, 7]
@@
In this work we are presenting band structure measurements on bulk 1T-VSe for different probe photon polarizations, which allows us to experimentally disentangle the contributions of the vanadium and selenium bands in the valence band structure.
We furthermore use the selective access to bands with different orbital character offered by pulses of different polarization.
@@
In Figs. 1a,b we show the layered crystal structure of bulk 1T-VSe and its hexagonal Brillouin zone (BZ).
@@ II.1 Photoelectron spectroscopy
The angle of incidence of the laser beam is 30 with respect to the surface normal.
The polarization of the 10.8 eV photons is changed by altering the polarization of the incoming beam in the last harmonic upconversion step, which is a third-harmonic generation stage performed in a cell filled with Xe gas. More details about the setup are reported in Ref. [17].
Due to the measurement geometry -polarized light always keeps a component parallel to the sample surface, whereas -polarized light probes only states with in plane components and cannot address states with a nodal plane parallel to the sample surface.
Tr-ARPES experiments were performed using the 800 nm fundamental beam of an amplified short pulse Ti:sapphire laser system as a pump beam and its 4th harmonic as a probe.
@@
Clean sample surfaces were prepared by cleaving in situ a bulk VSe crystal, purchased from HQ Graphene, using a kapton tape.
The background pressure of the ultra-high vacuum chamber, in which we were cleaving the samples amounted to mbar, while the measurements were performed at a background pressure of mbar.
The orientation of the sample on the sample holder was determined by investigating its low-energy electron diffraction pattern.
@@ II.2 Density functional theory calculations
Density functional theory calculations were performed using the Vienna ab initio package (VASP)[18, 19, 20] with the projector augmented wave method[21, 22]. Freestanding films were modeled with a 16 Åvacuum gap for the supercell. A plane-wave energy cutoff of 320 eV and a 16x16x10 k-mesh were adopted for the films. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) functional[23] was used; the bulk lattice constants are Åand Å. Spinorbit interactions were incorporated throughout the computations.
@@ III.1 Polarization dependent ARPES at equilibrium
The valence band structure of 1T-VSe originates from the V and the Se bands.
According to band structure calculations, these bands show different dispersion along the A direction of its hexagonal BZ.[24, 8, 25, 26]
Whereas four of those bands cross at the point of the BZ, at the A point these bands are separated from each other.
By performing ARPES using different photon energies and light polarizations, these bands are distinguishable by their different dispersion in the three dimensional momentum space and orbital band character.
DFT band structure calculations reported clear changes in the orbital character of the bands, despite strong hybridizations in certain regions of the BZ.[8, 24]
In the following we will compare ARPES measurements for different light polarizations to demonstrate that different parts of the valence band structure can be emphasized with - and -polarized light.
Figure 1c shows the valence band structure of bulk VSe, as probed at room temperature with photons of an energy of 10.8 eV.
Considering the dispersion of the valence bands and the inner potential of 16 eV reported by Falke et al. [25], we conclude that for free electron final states we are probing the sample band structure close to the A-point of the BZ.
We map the photoemission intensity as a function of binding energy with respect to the Fermi level () and electron momentum () along the AL and AH direction, respectively.
@@
Next we observe a shoulder indicating a third band with a larger effective mass than B.
Our experimental results for -polarized light are in qualitative agreement with previous ARPES studies at a photon energy of 11 eV presented by Falke et al.[25]
In the middle panel we repeated the same measurement with -polarized light and resolved the third band (C), that was barely visible with -polarized light.
Mapping the intensity difference between the two measurements, in the bottom panel we distinguish the bands with a significant out-of plane component (A and B) from the one with pure in-plane components (C).
Comparing our experimental results with our DFT band structure calculations, we can identify all three bands that are predicted to show up in the investigated part of the BZ.
@@
In agreement with calculations reported by Strocov et al. [15] and Majchrzak et al. [8] we attribute band A to the V band.
The orbital character of band B is controversial.
@@
According to Strocov et al. it corresponds to the Se state, whereas the band structure calculations reported by Majchrzak et al. show along AH two bands rather close to each other and strongly hybridizing leading to a more -like character close to the Fermi level and changing with increasing momentum.
@@
The dispersion and symmetry of band C along K agrees with the calculation reported by Strocov et al. for the Se.
@@
In our data we see a clear dependence on the polarization of the incident light, which leads to the conclusion that bands B and C have different band character, which supports our interpretation that band B is derived from the Se electrons and band C from the Se ones.
@@
Changing the probe photon energy to 6.2 eV, we move from the A plane to the plane of the BZ, if we consider the inner potential V of 16 eV as determined by Falke et al. [25] from low photon energy synchrotron measurements in the range from 9 to 17 eV. This inner potential is much larger than the value of 7.5 eV derived from synchrotron measurements at energies in the ranges of 65 to 160 eV (cf. Ref. [8], supporting material) and 845 to 960 eV (cf. Ref. [15]). The discrepancy between the two effective inner potential values for different incident photon energies could base on final state effects. Smaller variations of V can be observed in narrower photon energy ranges.[27] However, we assume that our photon energy of 6.2 eV is close enough to the photon energies used by Falke et al. and conclude that we probe very close to the point of the BZ.
@@
Note that this consideration does not include final state effects due to resonant excitation with our probe energy. The 6.2 eV photon energy reduces significantly our field of view in momentum space and binding energies. As we show in Fig. 2, we observe photoemission intensity in a binding energy range of eV. From these data we can deduce the sample work-function from the low energy cut-off () of the energy distribution curves (EDCs). We find eV, which is close to the value of eV reported by Claessen et al.[28] and confirms a clean and good quality surface. The photoemission intensity increases towards the Fermi-level. We observe a slight dispersion of the intensity maximum along the momentum axis . Overlaying our band structure calculations, we find that three of the four valence bands are positioned in the binding energy range where we observe enhanced photoemission intensity. Furthermore these three bands show comparable dispersion in our momentum space field of view. The fourth band instead is clearly separated from the region where we measure highest intensity. We mark the positions of the bands with black solid lines in the EDC shown at the right side of Fig. 2a, which is taken at the point (cf. integration window indicated by the dashed grey lines on top of the intensity distribution map). With the given energy resolution we cannot distinguish the different bands crossing in this energy and momentum region. Considering the polarization difference map we discussed above for 10.8 eV measurements, we obtain light blue color where the fourth band is predicted and red color in the range of the two uppermost bands that are strongly hybridizing (cf. Fig. 2b). The band with the highest binding energy shows Se character at the -point according to the calculations shown by Majchrzak et al.[8], whereas the other three are strongly hybridizing, so that the two -bands are getting significant -character. This is in contrast to the calculations published by Jolie et al.[24], but fits to our observation comparing different probe polarizations.
@@ III.2 Pump-probe spectroscopy
For investigating optically induced changes in the electron density of states in the vicinity to the Fermi level, we performed pump-probe experiments using 1.55 eV pump photon pulses with a fluence of mJ/cm.
According to our band structure calculations shown in Fig. 3a, there are selective values along M, AL and AH directions where resonant vertical transitions from bands with predominantly -like character (grey) to -like bands (black, blue) are predicted to be possible.
The unoccupied bands around 1 eV above the Fermi level are assigned to -like character in agreement with Ref. [24].
When pumping, we initially observe a decrease in the photoemission intensity below the Fermi level, indicating a depopulation of the bands (cf. the pump/non-pumped difference map shown in Fig. 3b) and an increase of the photoemission intensity just above the Fermi level.
@@
Indeed we are observing an almost symmetric shape of the difference EDC integrated in a small window around the -point at the delay of temporal overlap between pump and probe pulses (t, cf. Fig. 3b).
At later delays, however, the difference EDC gets highly asymmetric, clearly showing a residual enhanced intensity above the Fermi level and even a net increase of the signal below the Fermi level in comparison to equilibrium conditions (cf. Fig. 3b for the EDC difference curves at and 1 ps).
Tracing the evolution of the intensity distribution as a function of pump-probe delay at and at Å (see Fig. 3c), we observe an overall increase of the photoemission intensity in the range from -0.5 eV up to the Fermi level in the whole momentum range we are investigating, surviving for several picoseconds (cf. Fig. 3c).
At the right side of the intensity difference maps in Fig. 3c we show the difference EDCs averaged over all pump-probe delays larger than 1 ps, in order to display the spectral shape of the metastable state.
At we observe an overall intensity increase, which reaches its maximum around the Fermi level.
The signal seems to drop just below the Fermi level.
We attribute this drop to a depopulation of the vanadium and selenium bands, whose central positions are marked by the magenta lines.
This assignment is getting more clear when evaluating the difference EDCs at larger values.
The drop in intensity shifts here towards higher binding energies, according to the dispersion of the bands (cf. right panel of Fig. 3c).
In Fig. 3d we show the temporal evolution of the photoemission intensity for four distinct energy positions, indicated by the colored stripes in panel c.
The grey window is chosen in the uppermost part of our measurement window, well above the Fermi level, for extracting a transient that should be influenced mainly by our pump-probe cross correlation and hence allows us to deduce t for our dataset.
@@
In comparison to the black curve, all other curves show a delayed reach of their maximum intensity change.
@@
At Å the intensity change is suppressed to close to zero in the range of the upper three valence bands.
Majchrzak et al.[8] were reporting a depletion of the Se band reaching a plateau within the first 500 fs and not repopulating up to 5 ps after photoexcitation.
They were referring this to hot carriers affecting the low-energy soft phonon spectrum of the material and the electron-phonon coupling strength.
We conclude that the spectral shape of the difference EDC’s results from an intensity increase in the whole range from -0.5 eV up to the Fermi level, which is suppressed at some energies due to a long-lived depletion of the valence band population after photoexcitation.
In contrast to Majchrzak et al.,[8] we do not observe a net reduction of the photoemission intensity with respect to the equilibrium intensity due to a strong depopulation of the V band at the point that persists for several ps.
We refer this to the lower excitation fluence we are using in our experiments, which allows us to observe the net increase of the photoemission intensity.
@@
In Fig. 4 we compare the results at room temperature (well above T) discussed so far, with measurements performed at sample temperatures of 125 K (slightly above T) and 55 K (well below T).
Here we integrated over the whole momentum range from -0.1 to 0.1 Å.
@@
Cooling well below T (55 K) we find a more pronounced delayed band depletion at binding energies in the order of 0.5 eV.
After 1 ps, also in this case we find the same broad-range intensity increase as above T.
Comparing the difference EDCs for the three different temperatures (cf. Fig. 4b), we observe a clear downshift of the onset of the signal suppression due to valence band depletion with decreasing sample temperature.
@@
The optically induced intensity increase however does not seem to depend on the crystals phase.
@@
As shown in Fig. 4c, we evaluate the total electron yield (TEY) integrated in the complete field of view we are probing with 6.2 eV photons.
We observe an increase by 1 % of the equilibrium TEY upon 800 nm pumping.
@@
This observation is identical for starting in the normal and the CDW phase of the system and independent of the probe light polarization.
@@
Summing up our observations, we firmly exclude the observed intensity increase being related to a closure of the CDW gap due to optical excitation, since the signal shows no temperature dependence and our measurements at 300 K are far above the critical temperature of 110 K.
@@
We also exclude a shift of the entire band structure as being its origin, since 1.55 eV pump and HHG probe measurements published by Majchrzak et al. showed no indication for such a shift at higher excitation fluences.[8] Indeed, optically-induced band shifts should scale with the fluence.
@@
Furthermore, we demonstrated that our positive signal decreases following the equilibrium dispersion of the selenium bands (cf. Fig. 3c).
@@
At larger values the data of Majchrzak et al. show an indication of a drastic shift of the V band only, whose origin however is not discussed.
@@
Since our static data (cf. Fig. 1 and Fig. 2) clearly show that the V is probed with -polarized light only, we also exclude a shift in the vanadium band position to be the origin of our signal, since we can observe it likewise with both probe polarizations.
@@
A recent article claimed that a flat band appears in large parts of the bulk VSe BZ and can be shifted downwards via electron doping through donor adatoms adsorbed on the sample surface.[29]
@@
They speculated that polaron formation could be a possible origin of the discussed observation.
@@
An effect as electron doping via our pump pulses could be an explanation for our observation, however so far theoretical confirmation of flat bands or polarons in bulk VSe is lacking.
@@
Picosecond-scale modifications of the electronic density of states after optical excitation could also result from trapping electrons at defects in the crystal structure, as it was discussed e.g. for MoS monolayers.[30]

### v2

@@ Abstract
Complex materials encompassing different phases of matter can display new photoinduced metastable states differing from those attainable under equilibrium conditions. These states can be realized when energy is injected in the material following a non-equilibrium pathway, unbalancing the unperturbed energy landscape of the material. Guided by the fact that photoemission experiments allow for detailed insights in the electronic band structure of ordered systems, here we study bulk 1T-VSe2 in its metallic and charge-density-wave phase by time- and angle-resolved photoelectron spectroscopy. After near-infrared optical excitation, the system shows a net increase of the density of states in the energy range of the valence bands, in the vicinity of the Fermi level, lasting for several picoseconds. We discuss possible origins as band shifts or correlation effects on the basis of a band structure analysis. Our results uncover the possibility of altering the electronic band structure of bulk 1T-VSe2 for low excitation fluences, contributing to the understanding of light-induced electronic states.
@@
Layered transition-metal dichalcogenides (TMDCs) are attracting particular attention since they exhibit a variety of physical properties such as spin- and valley-coupled two-dimensional superconductivity, exciton Hall effect and exotic charge-density-wave (CDW) transitions.Manzeli et al. (2017)
@@
Path breaking works on light induced CDWs in compunds as VO2Morrison et al. (2014); Muraoka et al. (2018) and LaTe3 Kogar et al. (2020) triggered the search for similar effects in TMDCs.
@@
Photoexcitation can drive TMDC-CDW systems in metastable states surviving for several picoseconds, as reported for 1-TiSe2Hedayat et al. (2019); Duan et al. (2023), 1-TaSe2Shi et al. (2019) and 1-VSe2Majchrzak et al. (2021).
@@
Origins of these states are attributed to phonon bottleneck effects and optically-induced phase transitions from the CDW phase into a metallic phase.
@@
Duan et al. discovered that for TiSe2 an optically-induced metastable metallic phase can be realized far below the CDW transition temperature.
@@
This state, showing different properties than the equilibrium normal phase Duan et al. (2023), was connected to the optical triggering of a halted motion of the atoms via coherent electron-phonon coupling processes.
@@
In the case of 1T-VSe2, instead, phonon bottleneck effects after optical excitation in its metallic phase were discussed.Majchrzak et al. (2021)
@@
Recent interest has been drawn to the monolayer limit of VSe2 due to the enhanced CDW transition temperature and enlarged band gap.Biswas et al. (2021); Umemoto et al. (2019); Feng et al. (2018); Duvjir et al. (2018)
@@
Biswas et al.Biswas et al. (2021) report on ultrafast light induced insulator-metal transition in 2D VSe2.
@@
However, there are still open questions regarding the bulk properties of VSe2 to be answered, to gain a profound understanding of the physical mechanisms that determine its properties and phase transitions.
@@
With this work we contribute to the understanding of the optically-induced metastable state of 1T-VSe2 by performing time- and angle-resolved photoemission (tr-ARPES) experiments.
@@
We excite the material with 1.55 eV pump pulses and probe it at its point with 200 fs time and 50 meV energy resolution.
@@
Comparing results acquired at three different sample temperatures (well above, around and well below the CDW transition temperature of T110 KDiego et al. (2021)), we also address the influence of the initial phase of the system on the observed metastable modification.
@@
ARPES gives access to the valence band structure of a material, which in the case of 1T-VSe2 is given by the vanadium 3d (V3d) and selenium 4p (Se4p) bands.Terashima et al. (2003); Strocov et al. (2012); Majchrzak et al. (2021)
@@
Beyond the equilibrium band structure measurements, tr-ARPES allows studying the material response to an optical stimulus that drives the system far from equilibrium.
@@
This technique allows studying band-specific carrier dynamics which reveal e.g. information about the electron-electron and electron-phonon interactions in the system under investigation. Moreover, tr-ARPES was shown to be an adequate technique to study light-induced metastable states in the class of TMDC materials and beyond.Majchrzak et al. (2021); Fanfarillo et al. (2021); Duan et al. (2023); Hedayat et al. (2019); Shi et al. (2019)
@@
Here, we are presenting band structure measurements on bulk 1T-VSe2 for different probe polarizations, which allows us to experimentally disentangle the contributions of the vanadium and selenium bands in the valence band structure, exploiting the selective access to bands with different orbital symmetry as offered by a different probe polarization.
@@
In Fig.s 1(a),(b) we show the layered crystal structure of bulk 1T-VSe2 and its hexagonal Brillouin zone (BZ).
@@ II.1 Photoelectron spectroscopy
The angle of incidence of the laser beam is 30∘ with respect to the surface normal.
The polarization of the 10.8 eV photons is controlled by varying the polarization of the incoming beam in the last harmonic upconversion step, which is a third-harmonic generation stage performed in a cell filled with Xe gas. More details about the setup are reported in Ref. [Peli et al., 2020].
Due to the measurement geometry -polarized light always keeps a component parallel to the sample surface.
Tr-ARPES experiments were performed using the 800 nm (1.55 eV) fundamental beam of an amplified short pulse Ti:sapphire laser system as a pump beam and its 4th harmonic as a probe.
@@
Clean sample surfaces were prepared by cleaving in situ a bulk VSe2 crystal, purchased from HQ Graphene, using a kapton tape.
The background pressure of the ultra-high vacuum chamber, in which we were cleaving the samples was mbar, while the measurements were performed at a background pressure of mbar.
The required orientation of the samples was set via information from in-situ low-energy electron diffraction measurements.
@@ II.2 Density functional theory calculations
Density functional theory calculations were performed using the Vienna ab initio package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b) with the projector augmented wave methodBlöchl (1994); Kresse and Joubert (1999). Freestanding films were modeled with a 16 Å vacuum gap for the supercell. A plane-wave energy cutoff of 320 eV and a 16x16x10 -mesh were adopted for the films. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) functionalPerdew et al. (1996) was used; the bulk lattice constants are Å and Å. Spin-orbit interactions were incorporated throughout the computations. For joint-density of states calculations (cf. supplementary material) we use Wannier90 package.Pizzi et al. (2020)
@@ III.1 Polarization dependent ARPES at equilibrium
The valence band structure of 1T-VSe2 originates from the V3d and the Se4p bands.
According to band structure calculations, these bands show different dispersion along the A direction of its hexagonal BZ.Jolie et al. (2019); Majchrzak et al. (2021); Falke et al. (2021); Henke et al. (2020)
Whereas four of those bands overlap at the point of the BZ, at the A point they are separated from each other.
ARPES at different photon energies and light polarizations, makes these bands distinguishable by their different dispersion in the three dimensional momentum space and orbital band character.
DFT band structure calculations reported clear changes in the orbital character of the bands, despite strong hybridizations in certain regions of the BZ.Majchrzak et al. (2021); Jolie et al. (2019)
In the following we will compare ARPES measurements for different light polarizations to demonstrate that different parts of the valence band structure can be emphasized selectively with - and -polarized light.
Figure 1(c) shows the valence band structure of bulk VSe2, as probed at room temperature at 10.8 eV photon energy.
Considering the dispersion of the valence bands and the inner potential of 16 eV reported by Falke et al. Falke et al. (2021), we conclude that in the hypothesis of free-electron final states we are probing the sample band structure close to the A-point of the BZ.
We map the photoemission intensity as a function of binding energy with respect to the Fermi level () and electron momentum along the AL and AH direction (), respectively.
@@
Next we observe a shoulder suggesting the existence of a third band with a larger effective mass than B.
Our experimental results for -polarized light are in qualitative agreement with previous ARPES studies at a photon energy of 11 eV presented by Falke et al.Falke et al. (2021)
In the middle panel we repeat the same measurement with -polarized light unveilling a third band (C), that is barely visible with -polarized light.
Mapping the intensity difference between the two measurements, in the bottom panel we distinguish the bands with a significant out-of plane component or in plane nodal planes parallel to our analyzer slit () (A and B) from the one with in-plane components (C).Sugimoto et al. (2015); Moser (2017)
Comparing our experimental results with our DFT band structure calculations, we can identify the three bands predicted for the BZ sector we have investigated.
@@
In agreement with calculations reported by Strocov et al. Strocov et al. (2012) and Majchrzak et al. Majchrzak et al. (2021) we assign to the band A a V3d character.
The orbital character of band B is instead controversial.
@@
According to Strocov et al. it corresponds to the Se state, whereas the band structure calculations reported by Majchrzak et al. show two bands along AH, rather close to each other and strongly hybridizing, leading to a more -like character close to the Fermi level and changing with increasing momentum.
@@
The dispersion and symmetry of band C along K agrees with the calculation reported by Strocov et al. for the Se band.
@@
In our data we see a clear dependence on the polarization of the incident light, which leads to the conclusion that bands B and C have different origin.
From the orbital projection of the band structure calculations ( Fig.s S1 and S2 of the supplementary material), it becomes evident that band A is a pure V3d band, whereas band B and C have mixed character.
Both bands keep partial flavor and can be distinguished by their original atomic orbital symmetry.
The stronger-dispersing band B has imprinting, whereas band C is more .
This fits our experimental observation that B can be better probed with -polarized light.
@@
Changing the probe photon energy to 6.2 eV, we move from the AHL plane to the KM plane of the BZ, if we consider the inner potential V0 of 16 eV as determined by Falke et al. Falke et al. (2021) from low photon energy synchrotron measurements in the range from 9 to 17 eV. Instead, synchrotron measruements at photon energies above 65 eV report inner potentials of only 7.5 eV (cf. Ref. [Majchrzak et al., 2021], supporting material) and Ref. [Strocov et al., 2012]). The discrepancy between the two effective inner potential values for different incident photon energies could be based on final state effects. Smaller variations of V0 can be observed over narrower photon energy ranges.Tamura et al. (1985) However, we assume that our photon energy of 6.2 eV is sufficiently close to the photon energies used by Falke et al. and conclude that we probe the BZ very close to the point. Note that this consideration does not include final state effects originating from resonant excitation at our probe energy.
The 6.2 eV photon energy reduces significantly our field of view in momentum space and binding energies. As we show in Fig. 2, we observe photoemission intensity in a binding energy range of eV. From these data we can deduce the sample work-function from the low energy cut-off () of the energy distribution curves (EDCs). We find eV, which is close to the value of eV reported by Claessen et al.Claessen et al. (1990) and confirms a clean surface of good quality.
@@
The photoemission intensity increases towards the Fermi level. We observe a slight dispersion of the intensity maximum along . By comparing our band structure calculations with our experimental data, we find that three of the four valence bands are positioned in the binding energy range where we observe enhanced photoemission intensity. Instead, the fourth band is clearly separated from the region where we measure highest intensity. We mark the positions of the bands with black stripes in the EDC shown at the right side of Fig. 2(a), which is taken at the point (cf. integration window indicated by the dashed grey lines on top of the intensity distribution map). With the given energy resolution we cannot resolve the bands crossing in this energy and momentum region. For more clarity we report a polarization difference map in Fig. 2(b), which is telling us that we observe higher intensity with -polarized light (red) in the range of the upper three bands and slightly higher intensity with -polarized light (blue) in the binding energy range of the fourth band. The band with the highest binding energy has a Se4p character at the -point, according to the calculations reported by Majchrzak et al.Majchrzak et al. (2021), whereas the other three bands are strongly hybridizing, so that the two -bands have also a significant -character. This is in contrast to the calculations published by Jolie et al.Jolie et al. (2019), but fits to our observation when comparing different probe polarizations.
@@ III.2 Pump-probe spectroscopy
To investigate optically-induced changes in the electron density of states near the Fermi level, we performed pump-probe experiments using 1.55 eV pump photon pulses with an incident fluence of mJ/cm2.
According to our band structure calculations shown in Fig. 3(a), there are selective values along M, AL and AH directions where resonant vertical transitions from bands with predominantly -like character (grey) to -like bands (black, blue) are predicted to be possible.
The unoccupied bands around 1 eV above the Fermi level are assigned to -like character in agreement with Ref. [Jolie et al., 2019].
Out of equilibrium, we initially observe a decrease in the photoemission intensity below the Fermi level, indicating a depopulation of the bands (cf. the pump/non-pumped difference map shown in Fig. 3(b)) and an increase of the photoemission intensity just above the Fermi level.
@@
Indeed we are observing an almost symmetric shape of the difference EDC integrated in a small window around the -point at the delay of temporal overlap between pump and probe pulses (t0, cf. Fig. 3(b)).
At later delays, however, the difference EDC gets highly asymmetric, clearly showing a residual enhanced intensity above the Fermi level and even a net increase of the signal below the Fermi level in comparison to equilibrium conditions (cf. Fig. 3(b) for the EDC difference curves at and 1 ps).
Tracing the evolution of the intensity distribution as a function of pump-probe delay at and at Å-1 (see Fig. 3(c)), we observe an overall increase of the photoemission intensity in the range from -0.5 eV up to the Fermi level in the whole momentum range we are investigating, surviving for several picoseconds (cf. Fig. 3(c)).
At the right side of the intensity difference maps in Fig. 3(c) we report the difference EDCs averaged over all pump-probe delays larger than 1 ps, in order to display the spectral shape of the metastable state.
At we observe an overall intensity increase, which reaches its maximum around the Fermi level and then seems to drop just below the Fermi level.
We attribute this drop to a depopulation of the vanadium and selenium bands, whose central positions are marked as magenta stripes in Fig. 3(c).
This assignment is getting clearer when evaluating the difference EDCs at larger values.
The drop in intensity shifts here towards higher binding energies, according to the dispersion of the bands (cf. right panel of Fig. 3(c)).
In Fig. 3(d) we show the temporal evolution of the photoemission intensity for four distinct energies, indicated by the colored stripes in panel (c).
The grey window is chosen in the uppermost part of our measurement window, well above the Fermi level, to extract a transient that should be influenced mainly by our pump-probe cross correlation and hence allows us to deduce t0.
@@
In comparison to the grey curve, all other curves show a delayed reach of their maximum intensity change.
@@
At Å-1 the intensity change is suppressed to close to zero in the range of the upper three valence bands.
Majchrzak et al.Majchrzak et al. (2021) were reporting a depletion of the Se4p band reaching a plateau within the first 500 fs and not repopulating up to 5 ps after photoexcitation.
The authors were referring this to hot carriers affecting the low-energy soft phonon spectrum of the material and the electron-phonon coupling strength.
We conclude that the spectral shape of the difference EDCs results from an intensity increase in the whole range from -0.5 eV up to the Fermi level, which is suppressed at some energies due to a long-lived depletion of the valence band population after photoexcitation.
In contrast to Majchrzak et al.,Majchrzak et al. (2021) we do not observe a net reduction of the photoemission intensity with respect to the equilibrium intensity due to a strong depopulation of the V3d band at the point that persists for several ps.
We attribute this to the lower excitation fluence we are using in our experiments, which allows us to observe the net increase of the photoemission intensity.
@@
In Fig. 4 we compare the results at room temperature (well above Tc) discussed so far, with measurements performed at sample temperatures of 125 K (slightly above Tc) and 55 K (well below Tc).
Here we integrated over the whole momentum range from -0.1 to 0.1 Å-1.
@@
Cooling well below Tc (55 K) we find a more pronounced delayed band depletion at binding energies of the order of 0.5 eV.
After 1 ps, also in this case we find the same broad-range intensity increase as above Tc.
Comparing the difference EDCs for the three different temperatures (cf. Fig. 4(b)), we observe a clear downshift of the onset of the signal suppression due to valence band depletion with decreasing sample temperature.
@@
The optically-induced intensity increase however does not seem to depend on the crystals phase.
@@
As shown in Fig. 4(c), we evaluate the total electron yield (TEY) integrated in the complete field of view we are probing with 6.2 eV photons.
We observe an increase by 1 % of the equilibrium TEY upon 1.55 eV pumping.
@@
This observation is identical when starting from the normal or the CDW phase and does not depend on the probe polarization.
@@
Summing up our observations, we exclude that the observed intensity increase is related to a closure of the CDW gap due to optical excitation.
The signal shows no temperature dependence and our measurements at 300 K are far above the critical temperature of 110 K.
We also exclude the signal originating from a shift of the entire band structure.
@@
1.55 eV pump and HHG probe measurements published by Majchrzak et al. show no indication of such a shift at higher excitation fluences.Majchrzak et al. (2021) Since optically-induced band shifts should scale with the fluence, larger shifts are not expected at lower fluence.
@@
Furthermore, we demonstrated that our positive signal decreases following the equilibrium dispersion of the selenium bands (cf. Fig. 3(c)).
@@
At larger values the data of Majchrzak et al. show an indication of a drastic shift of the V3d band only, whose origin however is not discussed.
@@
Since our static data (cf. Fig. 1 and Fig. 2) clearly show that the V3d is probed with -polarized light only, we also exclude a shift in the vanadium band position to be the origin of our signal, since we can observe it likewise with both probe polarizations.
@@
A recent article claimed that a flat band is present in large parts of the bulk VSe2 BZ and can be shifted downwards via electron doping through donor adatoms adsorbed on the sample surface.Yilmaz and Vescovo (2022)
@@
These authors speculated that polaron formation could be at the origin of this mechanism.
Pump-induced electron doping could be an explanation for our observation.
@@
However, so far theoretical confirmation of flat bands or polarons in bulk VSe2 is lacking.
@@
Picosecond-scale modifications of the electronic density of states after optical excitation could also result from trapping electrons at defects in the crystal structure, as it was discussed e.g. for MoS2 monolayers.Wang et al. (2015)